I am extremely confused with how to solve fractions...especially fractions with two variables.
1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9
: I am extremely confused with how to solve fractions...especially fractions with two variables.
1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9

Answer by oscargut(666)

1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9 (-3)


1/2 x + 1/3 y = 1/3
-1/2 x -1 y = -2/3
adding both eq we have

-2/3 y = -1/3 then y=1/2
-1/2 x -1/2 =-2/3
-1/2 x = -2/3+1/2= -4/6 +3/6 =-1/6
x= 1/3
Solution: x=1/3 and y=1/2

Question 151455: I am extremely confused with how to solve fractions...especially fractions with two variables.
1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9
: I am extremely confused with how to solve fractions...especially fractions with two variables.
1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9

Answer by stanbon(18743)

1/2 x + 1/3 y = 1/3
1/6 x + 1/3 y = 2/9
-----------------------
Fractions are a nuisance; get rid of them.
Multiply thru 1st by 6
Multiply thru 2nd by 18
-------------------------
3x + 2y = 2
3x + 6y = 4
---------------
Subtract 1st from 2nd to get:
4y = 2
y = 1/2
-----------
Substitute into 3x+2y = 2 to solve for "x":
3x + 2(1/2) = 2
3x = 1
x = 1/3
---------------
Cheers,
Stan H.